Find the period of f(x) =cos(tanx+cotx) cos(tanx-cotx).

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
cos(tanx+cotx) cos(tanx-cotx) = (2/2) (cos(tanx+cotx)cos(tanx-cotx))
= (1/2) (cos (2tanx) + cos(2cotx))
{ using 2cosAcosB= cos(A+B) + cos(A-B) }
 
Period of f(x)= period of cos(2tanx) = period of cos(2cotx) = 2π/2 = π.
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)